A rectangle is 4 inches longer than it is wide. Its perimeter is 28 inches.
(a) Write a system of two equations in two variables that describes this information. Be sure to specify what each variable means.
(b) Solve your system graphically. Interpret your results.



Answer by mananth(16949)

length = x
width = y
x=y+4
x-y=4.............(1)
2x+2y=28
/2
x+y=14...........(2)

The lines intersect at (9,5)
width = 5
length = 9